Opened 11 years ago

Closed 11 years ago

Last modified 2 years ago

#57 closed flight processing (fixed)

CEH2007/06 (GB2007/12), flight day 192c/2007, Didcott

Reported by: mggr Owned by: anee
Priority: alpha 4 high Milestone: 2007 data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description

Data location: ~arsf/arsf_data/in_progress/2007/flight_data/uk/CEH2007_06-2007_192c_Didcott

Data arrived from ARSF via second SATA disk transfer at the start of Sept.

Scientific details:

  • ~/arsf_data/in_progress/2007/ARSF_Applications-GB_2007/GB07-12_Application_and_science_case.pdf

Only a single line of data.

Change History (18)

comment:1 Changed 11 years ago by mggr

  • Summary changed from GB2007/12 (and CEH2007/01?), flight day 240/2007, Buckingham to GB2007/12 (and CEH2007/06?), flight day 192c/2007, Didcott

Fixed summary line..

comment:2 Changed 11 years ago by mggr

  • Owner changed from mggr to anee

PI, Andrew Wilson, only wants raw data & processed applanix.

comment:3 Changed 11 years ago by anee

Combined operation unacceptably high – need to sort

comment:4 Changed 11 years ago by anee

location of base station (oxfr)
51 49 26.23733
-1 17 18.74360

location of 1st site
50 50 5.97049
-1 29 48.95154

location of 2nd site
50 45 30.30952
-2 26 47.89430

comment:5 Changed 11 years ago by mggr

Seems that the problem here was due to there being 3 sites visited on the one flight:

  • 192a (first site, Dorchester, LIDAR only)
  • 192b (second site, New Forest, LIDAR only)
  • 192c (third site, Didcott)

However, Didcott is actually only a single line of data so it doesn't look like another site - it's actually the straight part past the wiggle around 13:45 (time). The base station OXFR is the correct base station for this site. If you restrict the processing time range to just the part over Didcott (~13:35 - 13:55), it processes fine :)

comment:6 Changed 11 years ago by anee

Finished applanix processing using tips given above and a few small tweaks. Everything looks to be fine.

comment:7 Changed 11 years ago by mggr

Synced back into main repository.

comment:8 Changed 11 years ago by mggr

Sent on disk 3 to Andrew.

comment:9 Changed 11 years ago by mggr

  • Summary changed from GB2007/12 (and CEH2007/06?), flight day 192c/2007, Didcott to CEH2007/06 (GB2007/12), flight day 192c/2007, Didcott

Updating code per Gary's instructions:

GB07/12 was an ambiguous code that referred to eight different jobs, many of which were commissioned by CEH. To reduce confusion these were modified (in consultation with Andrew) to a CEH07 code. I know the log sheets still holds the code GB07/12 but I suggest that for the purposes of reporting these can be entered as 'CEH07/01 (GB07/12)'  for Buckingham (240), 'CEH07/06 (GB07/12)'  for Didcot (192b).  

comment:10 Changed 11 years ago by anee

  • Resolution set to fixed
  • Status changed from new to closed

comment:11 Changed 10 years ago by mggr

Archived at NEODC.

comment:12 Changed 2 years ago by dac

Hyperspectral

Processed data for this flight has been requested. Config file had already been generated.

Eagle header was named *_orig.hdr so renamed to match raw data.

comment:13 Changed 2 years ago by dac

Hyperspectral

Existing SCT values in config file were correct so processed using these. Level 3 data have some gaps in them - not sure why. Tried processing with different DEM and also get gaps. Created delivery but need to explain holes before delivery check.

comment:14 Changed 2 years ago by dac

Hyperspectral

Holes in mapped data are regions masked out due to incorrect smear correction. Have added to readme. Ready for delivery check.

comment:15 Changed 2 years ago by asm

Hyperspectral Delivery Check
Started

comment:16 Changed 2 years ago by asm

Hyperspectral Delivery Check

-> Project tidy is complaining about all names but everything seems fine.
-> Pixels masked due to smear correction catches the eye, note on readme makes sense and it is clear.
-> Everything else looks OK.

Zipping files and should be ready to send.

comment:17 Changed 2 years ago by asm

Hyperspectral Delivery Check

Mapped files have been zipped. Project ready to be delivered.

comment:18 Changed 2 years ago by dac

Hyperspectral files uploaded to FTP server (arsf7) on 2016-11-09.

Note: See TracTickets for help on using tickets.